OpenStudy (anonymous):

Find the product. (y 2)5 · y 8

OpenStudy (alekos):

Is that \[(y ^{2})^{5}.y ^{8}\]

OpenStudy (anonymous):

yes

OpenStudy (calculusxy):

You use the power rule which says: \[\large (x^a)^b = x^{a \times b} \] And then use the multiplication power rule with the same exponents: \[\large a^b \times a^c = a^{b+c} \]

OpenStudy (calculusxy):

Sorry i meant the same bases.

OpenStudy (calculusxy):

So you have: \[\large (y^2)^5 = y^{2 \times 5} = y^{10}\] And then: \[\large y^{10}\times y^8 = y^{10 + 8} = y^{18} \]